Grade 12th passPhysical Chemistry

Arrange electron proton deutron in increasing of specific charges

Ashwin Sarath
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Already ANswred this qstn. check your previous qstn.
Reposting it for the ease:
 
Since the charges would be same the specific charges would be for proton, electron and deutron will be => 
s1 = e/mp
s2 = e/me
s3 = e/2mp = e/2mp
mp is very very greater than mass of electron that is me.
 
Therfore the increasing order would be => 
s3 then s1 then me would be the order perhaps.